Find the points on the parabola y= x^2 closest to the point (0,1).
Answer by Alan3354(69443)

Find the points on the parabola y= x^2 closest to the point (0,1).
--------
the parabola is symmetrical about the y-axis and the point is on the y-axis.
---
Label the point in Q1 on the parabola (h,k)
---
d^2 = h^2 + (k-1)2
k = h^2
--> d^2 = k + k^2 - 2k + 1
d^2 = k^2 - k + 1
dk/dx = 2k - 1 = 0
k = 1/2
h = sqrt(2)/2
----------------
--> (sqrt(2)/2, 1/2)
& (-sqrt(2)/2, 1/2)